OpenStudy (anonymous):

A spacecraft is in orbit around a planet. The radius of the orbit is 2.9 times the radius of the planet (which is R = 71519 km). The gravitational field at the surface of the planet is 25 N/kg. What is the period of the spacecraft's orbit?

OpenStudy (anonymous):

\[T = \sqrt{\frac{ 4\Pi ^{2}r ^{3} }{ GM }}\]

OpenStudy (anonymous):

How do you do it if M is not given?

OpenStudy (anonymous):

can you use this formula \[M = \frac{ ar ^{2} }{ G }\]

OpenStudy (anonymous):

and what would a be?

OpenStudy (anonymous):

a = accleration due to gravity

OpenStudy (anonymous):

I can't seem to get the right answer. help please!

OpenStudy (anonymous):

put a=25 and try to solve

OpenStudy (anonymous):

I think r might be tripping me up. The radius of the orbit is 2.9 times the radius of the planet (which is R = 71519 km). but gravitational field is given at the surface of the earth

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that didn't give me the right answer either